OpenStudy (anonymous):

The sum of two numbers is 20. One number is three times as large as the second number. What are the two numbers?

OpenStudy (anonymous):

Which let statement would you use?

OpenStudy (anonymous):

x+3x=20

OpenStudy (anonymous):

x+3x=20 x=5 ,y=15

OpenStudy (anonymous):

A. Let x be the smaller number let 3x be the larger. B. Let x be the smaller number let 20x be the larger. C. Let x be the smaller number let x+3 be the larger.

OpenStudy (anonymous):

@Romero wt do u think

OpenStudy (anonymous):

x+3x=20

OpenStudy (anonymous):

So A) im a little confused between answer choice a and b

OpenStudy (anonymous):

x 3x which one is bigger

OpenStudy (anonymous):

3x

OpenStudy (anonymous):

"three times" what is times?

OpenStudy (anonymous):

addition or multiplication?

OpenStudy (anonymous):

multiplication

OpenStudy (anonymous):

Yeah so you cant use the last ones it has to be a

OpenStudy (anonymous):

@Romero thanks man I appreciate your helping me